a purse contains 100 coins that are nicks and dimes. The total value is $7. How many nickels and dimes are there?
Answer by checkley77(12844)

N+D=100 OR N=100-D
.05N+.10D=7.00
.05(100-D)+.10D=7.00
5-.05D+.10D=7.00
.05D=7-5
.05D=2
D=2/.05
D=40 NO. OF DIMES.
N+40=100
N=100-40
N=60 NO. OF NICKELS.
PROOF:
.05*60+.10*40*=7.00
3.00+4.00=7.00
7.00=7.00